Shift report 20030114 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 14 2003 - 15:52:03 EST

  • Next message: Hironori Ito: "Re: Shift report 20030113 00:00-08:00"
    ------------------------------------------------------------
    Date: Tue Jan 14 15:52:03 2003, Shift: 08:00-16:00
    Supervisor: AM, with: 
    Log book#13 pages 60 - 62
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam came about 10:15 a.m. and stayed
    there until ~ 2:30 p.m.
    2. Runs 7374 - 7378 have been taken.
    run 7375 was not written on HPSS, since it was
    just a test run for Dana.
    Ramiro worked on figuring out the background
    noise problem, rotating FS and BS (see the table
    in a log-book, page #61).
    The current settings are: both FS and BS are at
    8 degrees, MRS at 90 degrees. D5 magnet is on
    at 500A and in "B" polarity.
    3. DVM1 had problems. 
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Jan 14 2003 - 15:54:07 EST